What is the Slope and the Y-intercept?

- conc of dilu Trendline for conc of dilu R2 = 1 1.00 0.75 Concentration 0.50 0.25 0.00 0.2 0.3 0.4 0.5 0.6 0.7 od
Answer #1

Answer-

According to the given question-

conc of dilu Trendline for conc of dilu R2 = 1 1.00 0.75 Y2 Concentration 0.50 Slope = Y:-Y:/X2-X: 0.25 X1 X2 Intercept = c =

Here we have a graph and we have to find the slope as well as the intercept of the graph.

We know that the slope = Rise / Run = Y2 - Y1 / X2 - X1

Y2 = 0.50

Y1 = 0.25

X2 = 0.50

X1 = 0.25

Slope = 0.50 - 0.25 / 0.50 - 0.25 = 1

Slope = 1

While the intercept = c = point at which the line crosses Y-axis.

interceot = 0.125